OpenStudy (anonymous):

the rectangle below has a length of 60 and a width of 30. If the length is reduced by 50% and the width is increased by 50%, what will be the area of the rectangle?

OpenStudy (anonymous):

A=lw l=(60*.5) w=(30*1.5) A=30*45 A=

OpenStudy (anonymous):

@ChmE THANKS!